# Deep Generative Models for Detector Signature Simulation: An Analytical Taxonomy

### v1

@@ 6 Conclusion
In this paper, we offered an exhaustive and unified review of deep generative models for the efficient simulation of detector signatures in diverse experimental settings. As a result, we introduced a cohesive framework to articulate the task and provide an up-to-date and comprehensive taxonomy of algorithms. Subsequently, we conducted an in-depth survey of existing techniques and models, elaborating on their key characteristics. We then turned our attention to three primary application domains where deep generative models are making significant contributions to the simulation of detector signatures. In the end, we spotlighted the prevailing challenges in current research and offered a panoramic perspective on the future avenues for advancing deep generative models in the realm of efficient detector signature simulation.
@@ Acknowledgments
This research was part of Hosein Hashemi’s PhD thesis, supported by the collaborative project IDT-UM (Innovative Digitale Technologien zur Erforschung von Universum und Materie) funded by the German Federal Ministry of Education and Research (BMBF) and the Deutsche Forschungsgemeinschaft under Germany’s Excellence Strategy – EXC 2094 “ORIGINS” – 390783311. The authors would like to express their gratitude to the HEP ML Living Review [262], which provided invaluable assistance in compiling and locating an extensive body of research.

### v2

@@ 6 Conclusion
In this review, we offered an exhaustive and unified overview of deep generative models for the efficient simulation of detector signatures in diverse experimental settings. As a result, we introduced a cohesive framework to articulate the task and provide an up-to-date and comprehensive taxonomy of algorithms. Subsequently, we conducted an in-depth survey of existing techniques and models, elaborating on their key characteristics. We then turned our attention to three primary application domains where deep generative models are making significant contributions to the simulation of detector signatures. In the end, we spotlighted the prevailing challenges in current research and offered a panoramic perspective on the future avenues for advancing deep generative models in the realm of efficient detector signature simulation.
@@ Acknowledgments
This research was a chapter of Baran (Hosein) Hashemi’s PhD thesis [hashemi_deep_2024], supported by the collaborative project IDT-UM (Innovative Digitale Technologien zur Erforschung von Universum und Materie) funded by the German Federal Ministry of Education and Research (BMBF) and the Deutsche Forschungsgemeinschaft under Germany’s Excellence Strategy – EXC 2094 “ORIGINS” – 390783311. We thank Anna Zaborowska for discussions. The authors would like to express their gratitude to the HEP ML Living Review [hep_ml_community_living_nodate], which provided invaluable assistance in compiling and locating an extensive body of research.
[…]